you can ride your bicycle at 30 kph (km per hour) over a 30 km race comprising two 15 km laps. However, when you finish the first lap, your coach tells you that your average speed so far is only 15 kph. How fast must you ride to bring your average speed up to 30 kph.?
1
Expert's answer
2020-06-22T05:14:51-0400

Given, if "v_{av}=30\:kph" and suppose the required speed is "v" ,Thus


"v_{av}=30=\frac{30}{\frac{15}{15}+\frac{15}{v}}\\
\implies v\rightarrow \infty"

Thus, in practicality it is not possible to approach to average velocity 30 kph.